OpenStudy (anonymous):

Let u in R(n) be a nonzero row vector. (a) show that the n x n matrix A = u(t)u is symmetric and that rank(A)=1 (b)Show that the matrix P=I(n)+u(t)u is invertible.

OpenStudy (anonymous):

if u is just a row vector, then how is u(t)u anything but a single entry and the rank(A)=1, or do I just not understand what is going on here. @helder_edwin

OpenStudy (helder_edwin):

if u is a row vector then \[ \large u=(u_1\quad u_2\quad\dotsb\quad u_n) \] so \(u^t\cdot u\) is the matrix \[ \large A=\begin{pmatrix} u_1\\ u_2\\ \vdots\\ u_n \end{pmatrix}\cdot\begin{pmatrix} u_1 & u_2 & \dotsb & u_n \end{pmatrix} \] is a \(n\times n\) matrix

OpenStudy (helder_edwin):

symmetry is easy \[ \large A^t=(u^tu)^t=u^t(u^t)^t=u^tu=A \]

OpenStudy (helder_edwin):

i have to go. sorry. something urgent came up.

OpenStudy (anonymous):

I guess I was thinking about it as the row vector being first, then the column vector., I got symmetry that way, so then how is the rank =1

OpenStudy (anonymous):

okay, thank you.

OpenStudy (helder_edwin):

sorry i had to go. when u have A=BC then rank(A)<=rank(B) and rank(A)<=rank(C).

OpenStudy (helder_edwin):

this means that \[ \large \text{rank}(u^tu)\leq\text{rank}(u)=1 \] because u is a row vector.

OpenStudy (helder_edwin):

now since the only matrix with rank zero is the zero matrix then \[ \large \text{rank}(u^tu)>0 \] therefore rank(u^tu)=1.

OpenStudy (anonymous):

where did you get the rank(A)<=rank (B)? is that something that is just known to be true. A B and C are matrices?

OpenStudy (helder_edwin):

yes. it is a theorem of linear algebra.

OpenStudy (anonymous):

and the rank of u, would always 1 right, and since the zero matrix is the only rank 0, then rank(u(t)u) has to be 1.

OpenStudy (helder_edwin):

yes

OpenStudy (helder_edwin):

because \(u\neq0\)

OpenStudy (anonymous):

okay, so now for part (b), when I add the identity to A, I still get A right.. and since A is symmetric, then it has to be invertible too?

OpenStudy (helder_edwin):

no. u add 1 to every entry in the diagonal of A.

OpenStudy (anonymous):

ohh yes, right, and because its just the diagonal the only thing that changes, it would be invertible.

OpenStudy (helder_edwin):

i don't know how to prove that.

OpenStudy (helder_edwin):

perhaps using induction on the dimension n of the matrix????

OpenStudy (anonymous):

I thank you for your help and coming back to help with the problem. I will try to see if I can use induction.

OpenStudy (helder_edwin):

there is a theorem that says \[ \large \text{rank}(A+B)\leq\text{rank}(A)+\text{rank}(B) \]

OpenStudy (helder_edwin):

so rank(P)<=n+1

OpenStudy (helder_edwin):

if u use this u should have to disprove rank(P)=n+1 which is inmediat. since P is n*n. so rank(P)<=n

OpenStudy (helder_edwin):

now u would have to disprove rank(P)<n.

OpenStudy (anonymous):

there is a theorem in this section that I did not know how to apply... maybe it helps, i don't really know. its says.... nullity(A)+rank(A)=n

OpenStudy (helder_edwin):

right. i forgot that. so another strategy would be to prove that nullity(P)=0

OpenStudy (anonymous):

and why is the dim of the null space 0?

OpenStudy (helder_edwin):

is has to be. if so, then rank(P)=n and P would be invertible.

OpenStudy (anonymous):

okay, right b/c of that theorem, nullity(A) has to be zero, but if that is the case then why cant you just say its invertible b/c of the theorem?

OpenStudy (helder_edwin):

i mean. u should prove that nullity(P)=0. if u get this, then P would be invertible because rank(P) would equal n.

OpenStudy (helder_edwin):

i didn't express well my thoughts. sorry.

OpenStudy (anonymous):

I guess i am stuck on how to prove the dim=0

OpenStudy (helder_edwin):

if there is \(x\neq0\) such that Px=0 then \[ \large 0=Px=(I+u^tu)x=Ix+(u^tu)x=x+(u^tu)x \] so \[ \large (u^tu)x=-x \]

OpenStudy (anonymous):

It makes sense when I see it, I just can't figure out how to get things like this started. I appreciate your help though, I learn more for here than I do from the actual book.

OpenStudy (helder_edwin):

i can't see why this last equation should be wrong.

OpenStudy (helder_edwin):

sorry. i'm very tired.

OpenStudy (helder_edwin):

good luck with this problem.

OpenStudy (anonymous):

its okay, you have been a great help. i thank you again for everything.